A Bibliography of Graphics in TEX

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 2.13

Abstract

This bibliography records books and articles on the use of graphics in TEX documents.

Title word cross-reference

3 [Fin03, Fin04, Gon04, Roe97].

-dimensional [Fin03, Fin04].

158 [Moo98]. 1989 [Ano89]. 1992 [Zla92].

2-dimensional [Kri97]. 25th [SBH^{+}04]. 2D [Kot23].

3D [Kot23]. 3DLDF [Fin03, Fin04].

7th [Zla92].

’92 [Zla92].

ABC [Tob88b]. ACM [SIG77, SIG79]. ACM/SIGGRAPH [SIG77, SIG79]. Advanced [Her81]. Adventures [KM84]. algorithm [Ple99]. animations [Roe97]. Annual [SBH^{+}04]. ANSI [Ame85]. Applications [Cla90]. applied [BR89]. Approach [Har83, Har87, MB91]. Archeologia [Rah87a]. Asymptote [HBP04]. August [SBH^{+}04]. auspices [Rah87a]. AWK [Van86].

Barbara [USE97]. bars [Sch89]. Based [Nel85, HBP04]. biomedical [BR89]. bit [Tob85]. bitmaps [Föß91]. books [MB91]. Box [Van88]. boxed [Rec96]. bridge [vdL89]. Brush [Hob86]. \ [VR87].

C [SGN88]. California [USE97]. Calligraphy [Her67, Tob83]. can [vdL96b]. captions [Rec96]. Cartography [Her69]. Cemetery [Rah87a]. Center [And89]. challenge [Rog89]. charts [Kot23]. Chemical [HO87, NO90, Ram90, Fuj95]. chess [App88, Rub89]. child [vdL96b]. Chinese [HG83]. circuit [MB91]. Colophon [Ado]. Color [DRB89, Cor89]. Combining [Pic90a, Pic90b]. commands [Spr85]. Commercial [Nes87]. Committee [SIG77, SIG79, RH89]. Companion [GRM97, Rog97]. compatible [Pic90a]. comprehensive [BR89]. Computer [Ame85, Ame88, EKP84, FvD82, Har83, Har87, Her81, Her72, RS87, WH76, Knu86d, Rog89, Tob83]. Computers [Her67]. Conference [SBH^{+}04, USE97, Zla92, Spr85]. Congrès [Ano89]. Constraint [Nel85]. Constraint-Based [Nel85]. constructing [Wal88]. Contribution [WH76]. Cookbook [Ado85b]. Coordinates [WH76]. Creating [Roe97]. Creation [Wei92]. crosswords [Ham90]. cutting [SBL89]. Czechoslovakia [Zla92].

D [Gon04, Roe97]. DDA [Sal89]. degli [Rah87a]. dell’Arte [Rah87a]. Description [Ame87, DRB89]. Descriptors [DRB89]. Design [OB89, Rei88, Tob87]. Designing [Tob88a]. dessin [Hob01]. devices [Tob88a]. Diagram [Rey87]. diagrams [Jef91, Kot23, Lev90, MB91, Ohl95, Ram90]. Diffusion [Knu87a]. Digital [Kar87, Knu87a, SBH^{+}04, Uli87, Knu87b]. Digitized [Hob86]. dimensional [Fin03, Fin04, Kri97]. Directions [Knu79]. Discovering [Dam92]. Discussion [Spr85]. Display [Tuf83]. distributions [Car96]. do [vdL96b]. Document [Lam86, Lam94]. Documents [GRM97, Hoe90, Rec96, Dam92, Sch90, Wei92]. doggerel [Tob85]. Domain [USE97]. Domain-Specific [USE97]. Dot [Knu87a]. Draft [Ame88]. draw [Ham90, Kot23]. Drawing [BKW89, Hob95, KH97, Ohl95, Fin03, Fin04, Fuj95, Lev90, MB91, Ple99, Ren87, Wic88, Sch89]. Driver [Bee87, Nic91, RH89, Sch91]. DVI [RH89, Bee87, Sch91]. DVIPDF [Les97]. DVIPS [Arb86]. DVITOPS [Cla89].

editor [Wal88]. EEPIC [Kwo88]. effects [Tob88b]. Electronic [NO90]. empirical [Tob87]. Encoding [Cor89]. Enhancements [Pod86]. Environment [Nes87, Pod86, Kwo88, Sch89]. EPS [Rec96]. Erratum [Moo98]. Errors [Knu88]. European [Zla92]. EuroTEX [Zla92]. Extensions [Kwo88].

faces [Wuj88]. Family [Bee87, Tob84]. Feynman [Lev90, Ohl95]. Figures [BS90, Hoe90, Kne91, Rec96, Rec96]. File [CC88]. Floating [Rec96, Kne91]. Font [HG83]. Fonts [Hoe98, WH76, Win90, Knu87b, Tob84]. Format [CC88]. Formats [Kar87]. formatted [Wei92]. Formatter [HO87]. formulae [Ram90]. Formulas [HO87, Fuj95]. FORTRAN [Her69, Her70, BR89]. Foundations [RS87]. Frank [Rog97]. Functional [Ame87]. Fundamentals [FvD82]. Further [Wuj88].

generation [Win90]. Getting [Kar88]. GKS [Ame85, EKP84, HGDS83, HGDS86]. Glue [Van86]. GNUPLOT [Kot90, WKC^{+}90]. Goossens [Rog97]. GRAP [BK86]. graph [Hob01]. graphic [Nic91, SIG77, SIG79]. Graphical [Ame85, Ame88, HGDS83, HGDS86]. Graphics [Ame85, Ame88, Bee89, EKP84, FvD82, GRM97, Har83, Har87, Hoe98, Nel85, OB89, Rog97, WH76, And89, BR89, Ber90, CSB89, Cla92, Dam92, Ehr86, HBP04, Kar88, Ker82, Kot23, vdL96b, Lev90, Moo97, Moo98, Pic90a, Pic90b, Rec96, Rog89, Kri97, Sch90, Sow91, Wal88, Wei92, Wil89, Les97, Spr85, Car96, Ren89, vdL96a]. graphiques [Ano89]. Graphs [FSW99, Hob95, BK84, BK86, Ple99]. Greece [SBH^{+}04]. Group [SBH^{+}04]. Guide [Bee80, Lam86, Lam94, OQL88, Kot23]. GUTenberg [Ano89].

Halftone [Cla87]. Halftones [Knu87a, Knu87b]. Halftoning [Uli87, Cla91]. Hebrew [Pri81]. held [SBH^{+}04]. Hershey [WH76]. HI [SBL89]. HI-TEX [SBL89]. Hierarchical [Ame88]. High [Moo98, Van82, Kri97, Moo97]. High-Level [Van82]. histogram [Sch89]. HP3000 [Car81].

IBM [Pic90a]. II [Pic90b]. Illustrating [GRM97]. Image [CC88, Win90]. images [Kot23]. Implementation [OB89, Ber90]. included [Moo97, Moo98]. Including [Hei90, Sch90]. inclusion [CSB89]. incorporation [Wei92]. Independent [KM84, Ker81, Wil89]. Information [Ame85, Ame88, Tuf83]. integrating [Ren89]. Integration [Sow91, And89]. Interactive [Ame88, FvD82, WKC^{+}90]. interface [Nic91, Ren87]. International [SBH^{+}04]. Internazionale [Rah87a]. Introduction [Hob92, HGDS83, HGDS86, Jon89, Ros88]. issues [VR87]. Istituti [Rah87a]. IV [Her69].

Japanese [Sai87]. job [Rah87b]. jointly [SBH^{+}04]. Journal [Woo]. Juno [Nel85].

Kernel [Ame85, HGDS83, HGDS86].

Labelled [Jef91]. Labelling [Hoe90, Kar88]. labels [Moo97, Moo98, Wei92]. Language [Ado85a, Ado85b, BK86, KH97, OB89, Rei88, Van82, Woo, BK84, Gou86, HBP04, Ker82]. Languages [Ben86, DRB89, USE97]. laser [Pic90b]. LaserJet [Pic90a]. LaTEX [Sch89, Kot23]. LaTEX-paragraphs [Kne91]. left [KM87]. left-to-right [KM87]. letter [Pri81]. Level [Van82, Kri97]. Library [SGN88]. line [BR89, Ren87, Wil89]. line-drawing [Ren87]. Little [Ben86]. low [Tob88a]. low-res [Tob88a].

Machine [Wil89]. Machine-independent [Wil89]. Macintosh [Sch90]. Macros [Rey87, Wic88, Ham90, Van88]. Manual [AT&T83, Ado85a, Arb86, Cla89, Hob04, Lam86, Lam94, Nye88a, Nye88b, Arb88, Fin03, Fin04, Rey87, Wic87]. manuels [Hob01]. math [Rec96]. May [Ano89]. Meeting [SBH^{+}04]. Meta [HG83]. Meta-Font [HG83]. book [Knu86b]. METAPLOT [Wil89]. MetaPost [vdL96a]. Methodologies [Ren89]. Methods [Cla90, Sal89]. [Rub89, Tob87]. /MetaPost/PostScript [vdL96a]. Michel [Rog97]. Mittelbach [Rog97]. Mixing [KM87]. Modern [Knu86d]. Music [RS87, Gou86].

News [Ado]. nicely [BKW89]. Non [Sim90]. Non-traditional [Sim90].

observations [Tob87]. Occidental [WH76]. OCLC [Tob84]. October [USE97]. output [Cla87, Fin03, Fin04]. Overview [Ros88].

Packages [Gon04]. Page [DRB89]. paragraphs [Kne91]. Paris [Ano89]. Part [Car96, Rec96, Pic90b]. pasting [SBL89]. PC [Pic90a, Pic90b]. PC-compatible [Pic90a]. Pearls [Ben86]. PEX [Ros88]. PHIGS [Ame87, Bro85]. Physical [FSW99]. PIC [Ker82, Moo98, OB89, Moo97, Kri97]. Picture [KH97, Kwo88, OB89, Pod86, Sch89]. Picture-Drawing [KH97]. Pictures [Van82, Föß94, Hei90]. Pittsburgh [And89]. planar [Ple99]. Planning [SIG77, SIG79]. plots [Kot23, Van88]. Plotting [WKC^{+}90].

Portable [BS90, BR89, CSB89, Cla92]. PostScript [Ado85a, Ado85b, GRM97, Hol87, Rei88, Rot88]. pour [Hob01]. pp [Moo98]. Practical [Cla91]. practitioner [Kot23]. Prague [Zla92]. Preparation [Her70, Lam86, Lam94]. preparing [Ren89]. printers [Pic90b, Pic90a]. printing [Gou86, Rub89]. Problem [Bee89]. Proceedings [USE97, Zla92, SBH^{+}04]. Processing [NO90, Rah87b]. Program [Knu86a, Knu86c, Rei88, WKC^{+}90]. Programmer [Ame88, AT&T83]. Programming [Ben86, EKP84, Har83, Har87, Her69, Hol87, Nye88a]. Programs [Van86]. proposal [Ame88]. Protestant [Rah87a]. Protocol [SGN88]. Publication [Ado]. Publisher [Arb88]. Publishing [NO90]. Purpose [KH97].

quality [Moo97, Moo98]. Quantitative [Tuf83].

Rahtz [Rog97]. Real [Rot88]. reappraisal [vdL96a]. Reference [Ado85a, Lam86, Lam94, Nye88b, SGN88, Fin03, Fin04]. Repertoire [WH76]. Report [SIG77, SIG79, Sch91, RH89, Sai87]. Reports [Her70]. res [Tob88a]. research [BR89]. Review [Gon04]. Revision [Ame87, CC88]. right [KM87, Rah87b]. right-to-left [KM87]. Roma [Rah87a]. Roman [Tob84]. Rome [Rah87a]. routine [Lev90].

S [Nic91]. Santa [USE97]. scatter [Van88]. Scientific [Her72, BR89]. script [HBP04]. script-based [HBP04]. Sebastian [Rog97]. September [SBH^{+}04, Zla92]. SIGGRAPH [SIG77, SIG79]. Simulating [FSW99]. software [Nic91, Win90]. Some [Ham90, Tob87]. source [Pri81]. sous [Hob01]. Special [KH97, Tob88b, VR87]. Special-Purpose [KH97]. Specific [USE97]. Specification [CC88]. Specifying [Van82]. Standard [EKP84, Sch91, Cor89]. Standards [SIG77, SIG79, RH89]. State [Bee89]. statistical [Nic91, Ehr86]. Status [SIG77, SIG79]. Stereographic [Föß94]. Storia [Rah87a]. Strategies [Hoe98]. structural [Fuj95]. Structure [HO87, Ram90]. study [Rah87a]. Supercomputing [And89]. Symbols [WH76]. System [Ame85, Ame88, Her72, Her70, Hob97, HGDS83, HGDS86, Lam86, Lam94, Nel85, OQL88, BR89, Jon89, SGN88]. Systems [Ame85, Ame88, FSW99, Pic90a, Pic90b].

Tables [WH76]. Tag [CC88]. Techniques [WH76]. TEX [And89, OB89, Pic90b, Van88, Car81]. TEX-formatted [Wei92]. TEX/LaTEX [HO87]. TEXbook [Knu84]. TEX — [vdL96b, vdL96a, Spr85]. TEXhax [Spr85]. texpic} [OB89]. Text [HO87, NO90, MB91, Rah87b]. text-processing [Rah87b]. texts [KM87, KM87]. TEXtyl [Ren87]. TIFF [CC88]. TikZ [Kot23]. tools [Rah87b, Car96]. tour [Car96]. traditional [Sim90]. Trajectories [Hob86]. TransFig [BS90]. trees [BKW89, Epp85]. TROFF [KM84, Ker81]. TUG [SBH^{+}04]. tures [Wic88]. Turtle [vdL96b]. Tutorial [Ado85b, Kot90]. Type [WH76]. Typefaces [Kar87, Knu86d]. Typesetter [KM84, Ker81]. Typesetter-Independent [KM84, Ker81]. Typesetting [HO87, Knu79, WH76, BK84, BK86, Ker82, App88, vdL89]. Typographic [Her70]. Typography [Her81, Her72, Her69, SBH^{+}04].

Unbound [Hoe98]. Understanding [Bro85, Hol87]. undertaken [Rah87a]. Unione [Rah87a]. UNIX [AT&T83]. Use [Nes87]. User [Bee80, Cla89, Hob04, Lam86, Lam94, OQL88, Arb88, Rey87, Fin03, Fin04]. Users [SBH^{+}04]. Uses [Cla90, Sim90]. using [Föß94, Kot23, Moo97, Moo98, Rec96]. utilisateur [Hob01].

vector [HBP04]. Version [Kot90, Nye88a, Nye88b, OQL88]. Very [Kri97]. via [Rub89, vdL89]. Visual [Tuf83].

Window [Jon89, OQL88, SGN88]. World [Rot88].

x [Ram90, Föß91, Jon89, OQL88, SGN88]. x/y [Ram90]. X3.124 [Ame85]. X3.124-1985 [Ame85]. X3.144.1988 [Ame88]. Xanthi [SBH^{+}04]. Xerox [Cor89]. Xlib [Nye88b, Nye88a]. XML [SBH^{+}04]. Xy-pic [Moo98, Moo97, Kri97].

References

Adobe:colophon

[Ado] Adobe Systems Incorporated. Colophon—Adobe Systems News Publication.

Adobe:PLR85

[Ado85a] Adobe Systems Incorporated. PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10174-2. ix + 321 pp. LCCN QA76.73.P67 A33 1985.

Adobe:PLT85

[Ado85b] Adobe Systems Incorporated. PostScript Language Tutorial and Cookbook. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10179-3. x + 243 pp. LCCN QA76.73.P67 A34 1985.

ANSI:gks

[Ame85] American National Standards Institute, 1430 Broadway, New York, N. Y., 10018. Information Systems—Computer Graphics—Graphical Kernel System (GKS). ANSI X3.124-1985, 1985. Includes Fortran bindings to GKS.

ANSI:phigs+

[Ame87] American National Standards Institute, 1430 Broadway, New York, N. Y., 10018. PHIGS+ Functional Description, Revision 2.0, July 20 1987.

ANSI:phigs

[Ame88] American National Standards Institute, 1430 Broadway, New York, N. Y., 10018. Information Systems—Computer Graphics—Programmer’s Hierarchical Interactive Graphical System. Draft proposal X3.144.1988, 1988.

Andrews:TB10-2-177-178

[And89] Phil Andrews. Integration of TEX and graphics at the Pittsburgh Supercomputing Center. TUGboat, 10 (2):177–178, July 1989. ISSN 0896-3207.

Anonymous:TB10-1-118

[Ano89] Anonymous. GUTenberg Congrès: TEX et les graphiques, Paris, 16– 17 May 1989. TUGboat, 10(1):118, April 1989. ISSN 0896-3207.

Appelt:TB9-3-284-287

[App88] Wolfgang Appelt. Typesetting chess. TUGboat, 9(3):284–287, November 1988. ISSN 0896-3207.

Arbortext:1986

[Arb86] Arbortext Inc. Manual for DVIPS, 1986. See notes on LPEXT distributed with the software.

Publisher

[Arb88] Arbortext Inc. The Publisher User Manual. Ann Arbor, MI, USA, 1988.

ATT:UPM83-2

[AT&T83] AT&T. UNIX Programmer’s Manual, volume 2. Holt, Rinehart, and Winston, New York, NY, USA, 1983. ISBN 0-03-061743-X. vii + 616 pp. LCCN QA76.8.U65 B44 1983.

Beebe:plot79

[Bee80] Nelson H. F. Beebe. A user’s guide to <PLOT79>. Technical report, University of Utah, 1980.

Beebe:dvi-drivers

[Bee87] Nelson H. F. Beebe. A TEX DVI driver family. TEXniques, Publications for the TEX community, 5:71– 114, August 1987. Proceedings of the Eighth Annual Meeting of the TEX Users Group.

Beebe:tex-graphics

[Bee89] Nelson H. F. Beebe. TEX and graphics: The state of the problem. Cahiers GUTenberg, 2:13–53, 1989. ISSN 1140-9304. Presented to: Congrès GUTenberg, Paris, France, 16–17 May 1989.

Bentley:pic

[Ben86] Jon Louis Bentley. Programming pearls—little languages. Communications of the Association for Computing Machinery, 29(8):711–721, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Description of the pic language.

Berendt:TB11-2-190-194

[Ber90] Gerhard Berendt. On the implementation of graphics into TEX. TUGboat, 11(2):190–194, June 1990. ISSN 0896-3207.

Bentley/Kernighan:1984

[BK84] Jon Bentley and Brian W. Kernighan. grap—a language for typesetting graphs. Computing Science Technical Report 114, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1984. Also distributed as software documentation.

Bentley:grap

[BK86] Jon Louis Bentley and Brian W. Kernighan. GRAP—A language for typesetting graphs. Communications of the Association for Computing Machinery, 29(8):782–792, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruggemann-Klein:1989

[BKW89] Anne Brüggemann-Klein and Derrick Wood. Drawing trees nicely with TEX. Electronic Publishing —Origination, Dissemination, and Design, 3:101–115, 1989. CODEN EPODEU. ISSN 0894-3982.

Beebe:plot79-biomed

[BR89] Nelson H. F. Beebe and R. P. C. Rodgers. <PLOT79>: a comprehensive portable Fortran scientific line graphics system, as applied to biomedical research. Computers in Biology and Medicine, 19(6):385– 402, 1989.

Brown:UP85

[Bro85] Maxine Brown. Understanding PHIGS. Megatek Corporation, 9645 Scranton Rd, San Diego, CA 92121, USA, 1985. LCCN T385 .B761 1985.

Beck:TB11-3-373-380

[BS90] Micah Beck and A. Siegel. TransFig:

Portable figures for TEX. TUGboat, 11(3):373–380, 1990. ISSN 0896-3207.

Carnes:TB2-3-25

[Car81] Lance Carnes. TEX for the HP3000.

TUGboat, 2(3):25, November 1981. ISSN 0896-3207.

Carlisle:TB17-3-321

[Car96] David Carlisle. A LaTEX tour, Part 2: The tools and graphics distributions. TUGboat, 17(3):321–326, September 1996. ISSN 0896-3207.

Aldus:tiff

[CC88] Aldus Corporation and Microsoft Corporation. Tag image file format (TIFF) specification revision 5.0. Technical report, Aldus Corporation, 411 First Avenue South, Suite 200, Seattle, WA 98104, Tel:

(206) 622-5500, and Microsoft Corporation, 16011 NE 36th Way, Box 97017, Redmond, WA 98073-9717, Tel: (206) 882-8080, August 8 1988.

Clark:TB8-3-270

[Cla87] Adrian F. Clark. Halftone output from TEX. TUGboat, 8(3):270, November 1987. ISSN 0896-3207.

Clark:1989

[Cla89] James Clark. DVITOPS user manual. Unpublished machine-readable documentation., 1989.

Clark:TAU90

[Cla90] Malcolm Clark, editor. TEX Applications, Uses, Methods. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-912296-6.

Clark:TB12-1-157-165

[Cla91] Adrian Clark. Practical halftoning with TEX. TUGboat, 12(1):157–165, 1991. ISSN 0896-3207.

Clark:TB13-3-253

[Cla92] Malcolm Clark. Portable graphics in TEX. TUGboat, 13(3):253–260, October 1992. ISSN 0896-3207.

Xerox:color

[Cor89] Xerox Corporation. The Xerox Color Encoding Standard. Technical Report XNSS 288811, Xerox Systems Institute, March 1989.

Childs:TB10-1-44-46

[CSB89] Bart Childs, Alan Stolleis, and Don Berryman. A portable graphics inclusion. TUGboat, 10(1):44–46, April 1989. ISSN 0896-3207.

Damrau:TB13-3-315

[Dam92] Jackie Damrau. Discovering graphics in LaTEX documents. TUGboat, 13(3):315–321, October 1992. ISSN 0896-3207.

Dvonch:color-pdl

[DRB89] Mary Ann Dvonch, Paul Roetling, and Robert Buckley. Color descriptors in page description languages.

Proceedings of the Society for Information Display, 30(2):177–181, 1989. CODEN SIDPAA. ISSN 0734-1768, 0036-1496.

Ehrbar:TB7-3-171

[Ehr86] Hans Ehrbar. Statistical graphics with TEX. TUGboat, 7(3):171–175, October 1986. ISSN 0896-3207.

Enderle:CGP84

[EKP84] G. Enderle, K. Kansy, and G. Pfaff.

Computer Graphics Programming. GKS—The Graphics Standard. Symbolic Computation, Editor: J. Encarnação and P. Hayes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-11525-0, 3-540-11525-0. xvi + 542 pp. LCCN T385 .E531 1984.

Eppstein:TB6-1-31

[Epp85] David Eppstein. Trees in TEX.

TUGboat, 6(1):31–35, March 1985. ISSN 0896-3207.

Finston:2003:URM

[Fin03] Laurence D. Finston. 3DLDF user and reference manual: 3dimensional drawing with - output. ????, ????, 2003. ix

+ 269 pp. URL http://dante.
ctan.org/CTAN/graphics/3DLDF/

3DLDF.pdf. Manual edition 1.0 for 3DLDF version 1.0 September 2003.

Finston:2004:URM

[Fin04] Laurence D. Finston. 3DLDF user and reference manual: 3dimensional drawing with - output. ????, ????, 2004. x

+ 296 pp. URL http://dante.
ctan.org/CTAN/graphics/3DLDF/

3DLDF.pdf. Manual edition 1.1.5.1 for 3DLDF version 1.1.5.1 January 2004.

Fossmeier:TB12-2-229-232

[Föß91] Reinhard Fößmeier. X bitmaps in TEX. TUGboat, 12(2):229–232, 1991. ISSN 0896-3207.

Fossmeier:TB15-4-492

[Föß94] Reinhard Fößmeier. Stereographic pictures using TEX. TUGboat, 15 (4):492–495, December 1994. ISSN 0896-3207.

Frick:1999:SGP

[FSW99] Arne Frick, Georg Sander, and Kathleen Wang. Simulating graphs as physical systems. Dr. Dobb’s Journal of Software Tools, 24(8): 58, 60–64, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/.

Fujita:TB16-1-80

[Fuj95] Shinsaku Fujita. XΥMTEX for drawing chemical structural formulas. TUGboat, 16(1):80–88, March 1995. ISSN 0896-3207.

Foley:FIC82

[FvD82] James D. Foley and Andries van Dam. Fundamentals of Interactive Computer Graphics. The Systems Programming Series. Addison-Wes-ley, Reading, MA, USA, 1982. ISBN 0-201-14468-9. xx + 664 pp. LCCN T385 .F63 1982.

Goncalves:2004:FRM

[Gon04] Luis Nobre Gonçalves. - and a review of 3D packages. In Syropoulos et al. [SBH^{+}04], pages 112–124. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3

I58 2004. URL http://link.
springer-ny.com/link/service/
series/0558/tocs/t3130.htm;
http://www.springerlink.com/
openurl.asp?genre=issue&issn=
0302-9743&volume=3130;   http:
//www.springerlink.com/openurl.
asp?genre=volume&id=doi:10.1007/
b99374.

Gourlay:music-printing

[Gou86] John S. Gourlay. A language for music printing. Communications of the Association for Computing Machinery, 29(5):388–401, May 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goossens:1997:LGC

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The LaTEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp. LCCN Z253.4.L38G663 1997. US$39.75.

Hamilton-Kelly:TB11-1-103-119

[Ham90] Brian Hamilton Kelly. Some macros to draw crosswords. TUGboat, 11 (1):103–119, April 1990. ISSN 0896-3207.

Harrington:CGP83

[Har83] Steven Harrington. Computer Graphics—A Programming Approach. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-026751-0. xv + 448 pp. LCCN T385 .H34 1983. US$21.95.

Harrington:CGP87

[Har87] Steven Harrington. Computer Graphics—A Programming Approach. McGraw-Hill, New York, NY, USA, second edition, 1987. ISBN 0-07-026753-7. xiv + 466 pp. LCCN T385 .H34 1987. US$38.95.

Hammerlindl:2004:ASB

[HBP04] Andy Hammerlindl, John Bowman, and Tom Prince. Asymptote: a script-based vector graphics language. Faculty of Science, University of Alberta, Edmonton, AB, Canada, 2004. URL http:

                                        //asymptote.sourceforge.net/.

Asymptote is a powerful scriptbased vector graphics language for technical drawing, inspired by but with an improved C++-like syntax. Asymptote provides for figures the same highquality level of typesetting that LaTEX does for scientific text.

Heinz:1990

[Hei90] Alois Heinz. Including pictures in TEX. In Clark [Cla90]. ISBN 0-13-912296-6.

Hershey:calligraphy

[Her67] Allen V. Hershey. Calligraphy for computers. Technical Report TR-2101, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, August 1967.

Hershey:fortran-cartography

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969.

Hershey:fortran-typography

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:computer-typography

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1:373–385, 1972. CODEN CGIPBG. ISSN 0146-664X.

Hershey:1981:ACT

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Hobby:1983:CMF

[HG83] John Hobby and Gu Guoan. A Chinese Meta-Font. Report STAN-CS-83-974, Stanford University, Department of Computer Science, Stanford, CA, USA, 1983.

Hopgood:IGK83

[HGDS83] F. Robert A. Hopgood, Julian R.

Gallop, David A. Duce, and Dale C. Sutcliffe. Introduction to the Graphical Kernel System (GKS). Academic Press, New York, NY, USA, 1983. ISBN 0-12-355570-1. xi + 200 pp. LCCN T385 .I57 1983. A. P. I. C. Studies in Data Processing No. 19.

Hopgood:IGK86

[HGDS86] F. Robert A. Hopgood, Julian R. Gallop, David A. Duce, and Dale C. Sutcliffe. Introduction to the Graphical Kernel System (GKS). Academic Press, New York, NY, USA, second edition, 1986. ISBN 0-12-355571-X (paperback). xii + 250 pp. LCCN T385 .I571 1986. UK£12.50. Revised for the International Standard. A. P. I. C. Studies in Data Processing No. 28.

Haas/Kane:1987

[HO87] Roswitha T. Haas and Kevin C. O’Kane. Typesetting chemical structure formulas with the text formatter TEX/LaTEX. Computers and Chemistry, 11(4):251–271, 1987.

Hobby:1986:DBT

[Hob86] John Douglas Hobby. Digitized Brush Trajectories. Ph.D. dissertation, Department of Computer Science, Stanford University, Stanford, CA, USA, June 1986. 151

                                        pp. URL http://wwwlib.umi.com/
                                        dissertations/fullcit/8602484.

Also published as report STAN-CS-1070 (1985).

Hobby:1992:IM

[Hob92] John D. Hobby. Introduction to . In Zlatuška [Zla92], pages 21–36. ISBN 80-210-0480-0. Invited talk.

Hobby:1995:DGM

[Hob95] John D. Hobby. Drawing Graphs with . AT&T Bell Laboratories, Murray Hill, NJ, USA, 1995. 18 pp. URL http://ctan.

tug.org/tex-archive/macros/latex/
contrib/pdfslide/mpgraph.pdf.

Hobby:1997:MS

[Hob97] John D. Hobby. The System. ????, ????, December 30, 1997. 20 pp. URL file:/

//texlive-2004-11/texmf-dist/
doc/metapost/base/mpintro.pdf.

Hobby:2001:MDS

[Hob01] John D. Hobby. le dessin sous TEX: manuels de l’utilisateur pour et graph. Cahiers GUTenberg, 41(??): 175, 2001. ISSN 1140-9304. French translation by Pierre Fournier et Jean-Côme Charpentier.

Hobby:2004:UMM

[Hob04] John D. Hobby. A User’s Manual for . ????, ????, 2004. i + 85 pp. URL file:/

//texlive-2004-11/texmf-dist/
doc/metapost/base/mpman.pdf.

Hoenig:TB12-1-125-128

[Hoe90] Alan Hoenig. Labelling figures in TEX documents. TUGboat, 12(1): 125–128, 1990. ISSN 0896-3207.

Hoenig:1998:TUL

[Hoe98] Alan Hoenig. TEX Unbound:

LaTEX and TEX Strategies for Fonts, Graphics, & More. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-509686-X (paperback), 0-19-509685-1 (hardcover). ix + 580 pp. LCCN Z253.4.L38H64 1997. US$60.00 (hardcover), US$35.00 (paperback).

URL http://www.oup-usa.org/
gcdocs/gc_0195096851.html.

Holzgang:UPP87

[Hol87] David A. Holzgang. Understanding PostScript Programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1987. ISBN 0-89588-396-1. xxxii + 459 pp. LCCN QA76.73.P67 H65 1987.

Jeffrey:TB12-2-227-229

[Jef91] Alan Jeffrey. Labelled diagrams in . TUGboat, 12(2):227– 229, 1991. ISSN 0896-3207.

Jones:IXW89

[Jon89] Oliver Jones. Introduction to the X Window System. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-499997-5. xii + 511 pp. LCCN QA76.76.W56 J66 1989.

Karow:DFT87

[Kar87] Peter Karow. Digital Formats for Typefaces. URW-Verlag, Hamburg, Germany, 1987. ISBN 3-926515-01-5. 400 pp. LCCN Z253.3 .K371 1987.

Karney:1988

[Kar88] Charles Karney. Getting TEX labelling on graphics. Electronic mail to TEXhax, 24 May 1989., 1988.

Kernighan:ditroff

[Ker81] Brian W. Kernighan. A typesetterindependent TROFF. Technical Report Computer Science Report 91, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981.

Kernighan:pic

[Ker82] Brian W. Kernighan. PIC—A language for typesetting graphics. Software—Practice and Experience, 12 (1):1–22, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kamin:1997:SPL

[KH97] Samuel N. Kamin and David Hyatt. A special-purpose language for picture-drawing. In USENIX [USE97], pages 297– 310. ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http:/

/www.usenix.org/publications/
library/proceedings/dsl97/full_
papers/kamin/kamin_html/kamin.

html. This paper describes FPIC, an implementation in ML of a programmable extension to a PIC-like [Ker82] language. The authors raise interesting points about the difficulty of doing this correctly in ML, and conclude that a more powerful base language (e.g., Haskell) might be more suitable. FPIC includes support for both PostScript and LaTEX output.

Kahrs:ditroff

[KM84] Mark Kahrs and Lee Moore. Adventures with typesetter-independent TROFF. USENIX Summer Conference Proceedings, pages 258–269, June 12–15 1984. ISSN 1049-5606.

Knuth:TB8-1-14

[KM87] Donald Knuth and Pierre MacKay.

Mixing right-to-left texts with leftto-right texts. TUGboat, 8(1):14, April 1987. ISSN 0896-3207.

Kneser:TB12-1-28-30

[Kne91] Thomas Kneser. LaTEX-paragraphs floating around figures. TUGboat, 12(1):28–30, 1991. ISSN 0896-3207.

Knuth:1979:TMN

[Knu79] Donald E. Knuth. TEX and - : New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:1984:TB

[Knu84] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13447-0 (hardcover), 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$15.95 (paperback), US$32.95 (hardcover). Second printing, revised, October 1984. Sixth printing, revised, January 1986; also published as Computers & Typesetting, Vol. A. Twenty-fourth printing, summer 1993, contains final revisions. Twenty-first printing, June 1992.

Knuth:ct-b

[Knu86a] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986.

Knuth:ct-c

[Knu86b] Donald E. Knuth. The - book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:ct-d

[Knu86c] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:ct-e

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:halftone

[Knu87a] Donald E. Knuth. Digital halftones by dot diffusion. ACM Transactions on Graphics, 6(4):245–273, October 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knuth:TB8-2-135

[Knu87b] Donald E. Knuth. Fonts for digital halftones. TUGboat, 8(2):135–160, July 1987. ISSN 0896-3207.

Knuth:tex-errors

[Knu88] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988.

Kotz:gnuplot

[Kot90] David Kotz. GNUPLOT LaTEX Tutorial Version 2.0. Computer Science Department, Duke University, February 1990. See also [WKC^{+}90].

Kottwitz:2023:LGT

[Kot23] Stefan Kottwitz. LaTeX graphics with TikZ: a practitioner’s guide to draw 2D and 3D images, diagrams, charts, and plots in LaTEX using TikZ. Packt Publishing, Birmingham, UK, 2023. ISBN 1-80461-823-3 (paperback), 1-80461-704-0 (ebook). xviii + 279 pp. LCCN Z253.4.L38 K68 2023.

Rose:TB18-3-151

[Kri97] Kristoffer Høgsbro Rose. Very high level 2-dimensional graphics with TEX and Xy-pic. TUGboat, 18(3): 151–158, September 1997. ISSN 0896-3207.

Kwok:1988

[Kwo88] Conrad Kwok. EEPIC: Extensions to epic and LaTEX picture environment. Unpublished machinereadable document., 1988.

Lamport:1986:LDP

[Lam86] Leslie Lamport. LaTEX: a Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986.

Lamport:1994:LDP

[Lam94] Leslie Lamport. LaTEX: a Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-52983-1. xvi + 272 pp. LCCN Z253.4.L38L35

Lesenko:TB18-3-166

[Les97] Sergey Lesenko. DVIPDF and Graphics. TUGboat, 18(3):166–169, September 1997. ISSN 0896-3207.

Levine:CPC-58-181

[Lev90] Michael J. S. Levine. A LaTEX graphics routine for drawing Feynman diagrams. Computer Physics Communications, 58:181–198, 1990.

Maclenan:TB12-1-66-69

[MB91] M. P. Maclenan and G. M. Burns.

An approach to drawing circuit diagrams for text books. TUGboat, 12 (1):66–69, 1991. ISSN 0896-3207.

Moore:TB18-3-159

[Moo97] Ross Moore. High quality labels on included graphics, using Xy-pic. TUGboat, 18(3):159–165, September 1997. ISSN 0896-3207.

Moore:TB19-1-61

[Moo98] Ross Moore. Erratum: High quality labels on included graphics, using Xy-pic, TUGboat 1, no. 8(3), pp. 151–158. TUGboat, 19(1):61, March 1998. ISSN 0896-3207.

Nelson:1985:JCG

[Nel85] Greg Nelson. Juno, a constraintbased graphics system. Computer Graphics, 19(3):235–243, July 1985. CODEN CPGPBZ. ISBN 0-89791-166-0. ISSN 0097-8930.

Ness:tv-guide

[Nes87] David Ness. The use of TEX in a commercial environment. TEXniques, Publications for the TEX community, 5:115–123, August 1987. Proceedings of the Eighth Annual Meeting of the TEX Users Group.

Nicole:TB12-1-70

[Nic91] Olivier Nicole. A graphic driver to interface statistical software S and picTEX. TUGboat, 12(1):70–73, March 1991. ISSN 0896-3207.

Norris/Oakley:1990

[NO90] A. C. Norris and A. L. Oakley. Electronic publishing and chemical text processing. In Clark [Cla90]. ISBN 0-13-912296-6.

Nye:1988:XPM

[Nye88a] Adrian Nye. Xlib Programming Manual for Version 11, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-26-9, 0-937175-89-7. xxxiii + 615 pp. LCCN QA76.76.W56 D44 v.1

1988. URL http://www.oreilly.
com/catalog/9780937175262.

Nye:XRM88

[Nye88b] Adrian Nye. Xlib Reference Manual for Version 11, volume 2. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-27-7. xiv + 701 pp. LCCN QA76.76.W56 D44 v.2 1988.

Olejniczak-Burkert:TB10-4-627-637

[OB89] Rolf Olejniczak-Burkert. texpic— Design and Implementation of a Picture Graphics Language in TEX à la pic. TUGboat, 10(4):627–637, December 1989. ISSN 0896-3207.

Ohl:1995:DFD

[Ohl95] Thorsten Ohl. Drawing Feynman diagrams with LaTEX and - . Computer Physics Communications, 90(2–3):340–354, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Oreilly:XWS88

[OQL88] Tim O’Reilly, Valerie Quercia, and Linda Lamb. X Window System User’s Guide for Version 11, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-29-3.

Pickrell:TB11-1-26-31

[Pic90a] Lee S. Pickrell. Combining graphics with TEX on IBM PC-compatible systems and LaserJet printers. TUGboat, 11(1):26–31, April 1990. ISSN 0896-3207.

Pickrell:TB11-2-200-206

[Pic90b] Lee S. Pickrell. Combining graphics with TEX on PC systems with laser printers, part II. TUGboat, 11 (2):200–206, June 1990. ISSN 0896-3207.

Plestenjak:1999:ADP

[Ple99] Bor Plestenjak. An algorithm for drawing planar graphs. Software —Practice and Experience, 29(11): 973–984, September 1999. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=
63501203;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?ID=
                                        63501203&PLACEBO=IE.pdf.

epic

[Pod86] Sunil Podar. Enhancements to the picture environment of LaTEX. Technical Report 86-17, Department of Computer Science, S.U.N.Y, 1986. Version 1.2: July 14, 1986.

Price:TB2-1-122

[Pri81] Lynne A. Price. Hebrew letter (with source). TUGboat, 2(1):122, February 1981. ISSN 0896-3207.

Rahtz:1987b

[Rah87a] Sebastian Rahtz. The Protestant Cemetery, Rome: a study undertaken under the auspices of the Unione Internazionale degli Istituti di Archeologia Storia e Storia dell’Arte in Roma. Opuscula Romana, 16:149–167, 1987.

Rahtz:1987a

[Rah87b] Sebastian Rahtz. The right textprocessing tools for the job? CS-MAN 87-4, Computer Science, University of Southampton, 1987.

Ramek:1990:CSF

[Ram90] Michael Ramek. Chemical structure formulae and x/y diagrams with TEX. In Clark [Cla90]. ISBN 0-13-912296-6.

Reckdahl:1996:UAG

[Rec96] Keith Reckdahl. Using EPS graphics in LaTEX2ε documents, Part 2: Floating figures, boxed figures, captions, and math in figures. TUGboat, 17(3):288–310, September 1996. ISSN 0896-3207.

Reid:1988:PLP

[Rei88] Glenn C. Reid. PostScript Language Program Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-14396-8. xii + 224 pp. LCCN QA76.73.P67 R45 1988. US$22.95.

URL http://www.rightbrain.
com/rightbrain.shtml.

Renner:textyl

[Ren87] John S. Renner. TEXtyl: a linedrawing interface for TEX. Technical Report OSU-CISRC-4/87-TR9, Department of Computer Science, Ohio State University, March 1987.

Renfrow:TB10-4-607

[Ren89] J. T. Renfrow. Methodologies for preparing and integrating Post-Script graphics. TUGboat, 10(4): 607–626, December 1989. ISSN 0896-3207.

Reynolds:1987

[Rey87] J. C. Reynolds. User’s manual for diagram macros. Unpublished machine-readable document., December 1987.

Reid:TB10-2-188-191

[RH89] Tom Reid and Don Hosek. Report from the —DVI— driver standards committee. TUGboat, 10(2):188– 191, July 1989. ISSN 0896-3207.

Roegel:1997:CAM

[Roe97] Denis Roegel. Creating 3D animations with . TUGboat, 18(4):274–283, December 1997. ISSN 0896-3207. URL http://

ctan.tug.org/tex-archive/graphics/
metapost/contrib/macros/3d/doc/
paper1997corrected.pdf.

Rogers:TB10-1-39-44

[Rog89] David F. Rogers. Computer graphics and TEX—a challenge. TUGboat, 10(1):39–44, April 1989. ISSN 0896-3207.

Rogers:TB18-4-246

[Rog97] David F. Rogers. The LaTEX Graphics Companion, by Michel Goossens, Sebastian Rahtz and Frank Mittelbach. TUGboat, 18(4):246, December 1997. ISSN 0896-3207.

Rost:pex

[Ros88] Randi J. Rost. PEX introduction and overview. Technical Report Version 3.20, Digital Equipment Corporation, Workstation Systems Engineering, April 1988. This document is present in the X Window System Version 11 Release 3 in

the file X11/X11/doc/extensions/
pex/doc/intro/doc.ms.

Roth:RWP88

[Rot88] Stephen E. Roth, editor. Real World PostScript. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-06663-7. xiv + 383 pp. LCCN Z286.D47 R4 1988.

Roads:FCM87

[RS87] Curtis Roads and John Strawn, editors. Foundations of Computer Music. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-18114-2. xiii + 712 pp. LCCN ML1092 .F7 1985.

Rubinstein:TB10-2-170-172

[Rub89] Zalman Rubinstein. Chess printing via and TEX. TUGboat, 10(2):170–172, July 1989. ISSN 0896-3207.

Saito:TB8-2-103

[Sai87] Yasuki Saito. Report on JTEX: A Japanese TEX. TUGboat, 8(2):103, July 1987. ISSN 0896-3207.

Salomon:TB10-2-207-216

[Sal89] David Salomon. DDA methods in TEX. TUGboat, 10(2):207–216, July 1989. ISSN 0896-3207.

Syropoulos:2004:TXD

[SBH^{+}04] Apostolos Syropoulos, Karl Berry, Yannis Haralambous, Baden Hughes, Steven Peter, and John Plaice, editors. TEX, XML, and Digital Typography: International Conference on TEX, XML, and Digital Typography, held jointly with the 25th Annual Meeting of the TEX Users Group, TUG 2004, Xanthi, Greece, August 30–September 3, 2004: Proceedings, volume 3130 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3

I58 2004. URL http://link.
springer-ny.com/link/service/
series/0558/tocs/t3130.htm;
http://www.springerlink.com/
openurl.asp?genre=issue&issn=
0302-9743&volume=3130;   http:
//www.springerlink.com/openurl.
asp?genre=volume&id=doi:10.1007/
b99374.

Spivak:TB10-2-164-165

[SBL89] Michael Spivak, Michael Ballantyne, and Yoke Lee. HI-TEX cutting & pasting. TUGboat, 10(2):164– 165, July 1989. ISSN 0896-3207.

Schopf:TB10-1-105-107

[Sch89] Rainer Schöpf. Drawing histogram bars inside the LaTEX picture– environment. TUGboat, 10(1):105– 107, April 1989. ISSN 0896-3207.

Schwer:TB11-2-194

[Sch90] Len Schwer. Including Macintosh graphics in LaTEX documents. TUGboat, 11(2):194–200, June 1990. ISSN 0896-3207.

Schrod:TB12-2-232-233

[Sch91] Joachim Schrod. Report on the DVI driver standard. TUGboat, 12(2): 232–233, 1991. ISSN 0896-3207.

Scheifler:XWS88

[SGN88] Robert W. Scheifler, James Gettys, and Ron Newman. X Window System: C Library and Protocol Reference. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1988. ISBN 1-55558-012-2. xxix + 701 pp. LCCN QA76.76.W56 S34 1988.

SIGGRAPH:core77

[SIG77] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH— Computer Graphics, 11(3), 1977.

SIGGRAPH:core79

[SIG79] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH— Computer Graphics, 13(3), August

Simpson:1990:NTU

[Sim90] Rick Simpson. Non-traditional uses of . In Clark [Cla90]. ISBN 0-13-912296-6.

Sowa:1991:IGT

[Sow91] Friedrich Sowa. Integration of graphics into TEX. TUGboat, 12(1): 58–63, 1991. ISSN 0896-3207.

Spragens:TB6-2-66

[Spr85] Alan Spragens. Graphics commands for TEX — Discussion in TEXhax conference. TUGboat, 6(2):66–68, July 1985. ISSN 0896-3207.

Tobin:TB4-1-26

[Tob83] Georgia K. M. Tobin. Computer calligraphy. TUGboat, 4(1):26, April 1983. ISSN 0896-3207.

Tobin:TB5-1-36

[Tob84] Georgia K. M. Tobin. The OCLC Roman family of fonts. TUGboat, 5 (1):36, May 1984. ISSN 0896-3207.

Tobin:TB6-1-12

[Tob85] G. K. M. Tobin. A bit of doggerel. TUGboat, 6(1):12, March 1985. ISSN 0896-3207.

Tobin:TB8-1-26

[Tob87] Georgia K. M. Tobin. Some empirical observations on design. TUGboat, 8(1):26, April 1987. ISSN 0896-3207.

Tobin:TB9-2-126-128

[Tob88a] Georgia K. M. Tobin. Designing for low-res devices. TUGboat, 9(2):126– 128, August 1988. ISSN 0896-3207.

Tobin:TB9-1-15-18

[Tob88b] Georgia K. M. Tobin. The ABC’s of special effects. TUGboat, 9(1):15– 18, April 1988. ISSN 0896-3207.

Tufte:VDQ83

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X. 197 pp. LCCN K27.S8 T84. US$34.00.

Ulichney:DH87

[Uli87] Robert Ulichney. Digital Halftoning. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-21009-6. xiv + 362 pp. LCCN T385 .U451 1987.

USENIX:1997:PCD

[USE97] USENIX, editor. Proceedings of the Conference on Domain-Specific Languages, October 15–17, 1997, Santa Barbara, California. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997.

VanWyk:ideal

[Van82] Christopher J. Van Wyk. A highlevel language for specifying pictures. ACM Transactions on Graphics, 1(2):163–182, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanWyk:awk

[Van86] Christopher J. Van Wyk. AWK as glue for programs. Software— Practice and Experience, 16(4):369– 388, April 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

vanHaagen:TB9-2-189-192

[Van88] A. J. Van Haagen. Box plots and scatter plots with TEX macros. TUGboat, 9(2):189–192, August 1988. ISSN 0896-3207.

vanderLaan:TB10-1-113-116

[vdL89] C. G. van der Laan. Typesetting bridge via LaTEX. TUGboat, 10(1): 113–116, April 1989. ISSN 0896-3207.

Laan:TB17-3-269

[vdL96a] Kees van der Laan. Graphics and TEX — A reappraisal of /MetaPost/Post-Script. TUGboat, 17(3):269–279, September 1996. ISSN 0896-3207.

Laan:TB17-2-222

[vdL96b] Kees van der Laan. Turtle graphics and TEX — a child can do it. TUGboat, 17(2):222–228, June 1996. ISSN 0896-3207.

Vanderburg:TB8-3-291-300

[VR87] Glenn L. Vanderburg and Thomas J.

Reid. \special issues. TUGboat, 8 (3):291–300, November 1987. ISSN 0896-3207.

Waldschmidt:1988

[Wal88] H. Waldschmidt. An editor for constructing graphics with TEX. Computer Graphics Forum, 6(4):359– 364, 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:TB13-3-330

[Wei92] Neil A. Weiss. Creation and incorporation of PostScript graphics with TEX-formatted labels into TEX documents. TUGboat, 13(3): 330–334, October 1992. ISSN 0896-3207.

Wolcott:1976:CCT

[WH76] Norman M. Wolcott and Joseph Hilsenrath. A contribution to computer typesetting techniques. Tables of coordinates for Hershey’s repertoire of occidental type fonts and graphics symbols. Technical Report PB-251 845, U. S. National Bureau of Standards, April 1976. NBS Special Publication 424.

Wichura:PM87

[Wic87] Michael J. Wichura. The picTEX manual. Number 6 in TEXniques: publications for the TEX community. TEX Users Group, Providence, RI, USA, 1987.

Wichura:TB9-2-193-197

[Wic88] Michael J. Wichura. picTEX: Macros for drawing pictures. TUGboat, 9(2):193–197, August 1988. ISSN 0896-3207.

Wilcox:TB10-2-179-187

[Wil89] Patricia Wilcox. METAPLOT: Machine-independent line graphics for TEX. TUGboat, 10(2):179–187, July 1989. ISSN 0896-3207.

Winckler:1990:TFI

[Win90] Jörg Winckler. TEX fonts in image generation software. In Clark [Cla90]. ISBN 0-13-912296-6.

Williams:gnuplot

[WKC^{+}90] Thomas Williams, Colin Kelley, John Campbell, David Kotz, and Russell Lang. GNUPLOT—An Interactive Plotting Program, 31 August 1990. Available in several Internet archives, including the Free Software Foundation collection on

prep.ai.mit.edu. GNUPLOT can

produce output for many different devices, including LaTEX picture mode, PostScript, and the X Window System. See also [Kot90].

Wood:plj

[Woo] Patrick Wood, editor. Post-ScriptLanguage Journal. Pipeline Associates, Inc., P. O. Box 5763, Parsippany, NJ 07054.

Wujastyk:TB9-3-246-251

[Wuj88] Dominik Wujastyk. Further faces.

TUGboat, 9(3):246–251, November 1988. ISSN 0896-3207.

Zlatuska:1992:EPE

[Zla92] Jiří Zlatuška, editor. EuroTEX ’92: Proceedings of the 7th European TEX Conference, Prague, Czechoslovakia, September 14–18, 1992, Proceedings of the European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-